OpenStudy (anonymous):

in some year election for president there was 84.4 million voters,where as in the next year was 126.2 million voters. Find the percent change in the number of voters

OpenStudy (anonymous):

126.2=84.4*(1+p/100) 1+p/100=126.2/84.4 p/100=0.495 p=49.5